# Consistency of EFT illuminated via relative entropy: A case study in scalar field theory

### v1

@@ 6 Summary and outlook
We revisited an EFT of single-field inflation as an application of this relative entropy method. The experimentally allowed regions of the non-linear parameter in EFT (65) were shown to be limited potentially when the EFT description is valid in low-energy regions. The relative entropy of the two scalar field theories (one of the UV theories of the EFT of single-field inflation) was evaluated in detail by analytical and numerical calculations for two typical cases: the U(1) symmetry and the U(1) symmetry-breaking theory. In the perturbative calculations, it is found that the U(1) symmetry theory respects the non-negativity of relative entropy, while the U(1) symmetry-breaking theories may not. Although the conventional method of perturbative unitarity based on scattering amplitude has the advantage of constraining the absolute values of the non-linear parameters (in two scalar fields theory), our study showed that a novel advantage, namely a severe constraint on its sign, can be exposed by shedding new light on the relative entropy. In addition, we presented the regions of the non-linear parameter where the EFT description is valid in a UV theory-independent way, which would provide valuable insights into identifying models describing physical phenomena complementary to the experimental data.
We will mention a couple of things to the prospects of this work at the end. First, constraints on the EFT of single-field inflation might be generalized to EFTs of multi-field inflation, which have also been studied in the context of the positivity bounds Freytsis:2022aho . Second, although our work has in mind the particular class of UV theories presented in Sec. 2, it would be interesting to discuss possible extensions to UV theory independent arguments (more specifically, an approach based on calculations of the relative entropy from the EFT). Third, our relative entropy perspective can be applied to other EFTs describing Lorentz symmetry-breaking phenomena, such as superfluids. Since conventional positivity bounds rely on the Lorentz symmetry, we expect our relative entropy perspective, which does not depend on Lorentz symmetry, to produce more diverse physics results.
@@ Acknowledgements
We are grateful to Toshifumi Noumi for his helpful comments and the opportunity to initiate this work. DU would like to thank the CERN Theory Department for their financial support and hospitality during this work.
@@ Appendix A Field redefinition of heavy fields
Appendix B Higher order of perturbation expansion in the effective action
@@
    U(1) symmetric potential — Upon substituting eq. (76) into eq. (135), we obtain the Euclidean effective actions, up to the sixth order of $\lambda$, as
@@
    U(1) symmetry-breaking potential — Upon substituting eq. (86) into eq. (135), we obtain the Euclidean effective actions, up to the sixth order of $\lambda$, as
@@
Appendix C Relation between mass bases and interaction bases of the fields

### v5

@@ 5 Summary and outlook
We revisited an EFT of single-field inflation as an application of this relative entropy method. The experimentally allowed regions of the non-linear parameter in EFT (66) were shown to be limited potentially when the EFT description is valid in low-energy regions. The relative entropy of the two scalar field theories (one of the UV theories of the EFT of single-field inflation) was evaluated in detail by analytical and numerical calculations for two typical cases: the U(1) symmetry and the modified theory. In the perturbative calculations, it is found that the U(1) symmetry theory in subsubsection 3.1.1 respects the non-negativity of relative entropy, while the modified theories may not. Although the conventional method of perturbative unitarity based on scattering amplitude has the advantage of constraining the absolute values of the non-linear parameters (in two scalar fields theory), our study showed that a novel advantage, namely a severe constraint on its sign, can be exposed by shedding new light on the relative entropy. In addition, we presented the regions of the non-linear parameter where the EFT description is valid in a UV theory-independent way, which would provide valuable insights into identifying models describing physical phenomena complementary to the experimental data.
We will mention a couple of things to the prospects of this work at the end. First, constraints on the EFT of single-field inflation might be generalized to EFTs of multi-field inflation, which have also been studied in the context of the positivity bounds Freytsis:2022aho . Second, although our work has in mind the particular class of UV theories presented in Sec. 2, it would be interesting to discuss possible extensions to UV theory independent arguments (more specifically, an approach based on calculations of the relative entropy from the EFT). Third, a natural extension of this work would be to generalize it to open systems. In cosmology, non-unitary time evolution may play a significant role Brahma:2022yxu ; Brahma:2024yor , and its effects on EFT could be explored by considering relative entropy. Forth, our relative entropy perspective can be applied to other EFTs describing Lorentz symmetry-breaking phenomena, such as superfluids. Since conventional positivity bounds rely on the Lorentz symmetry, we expect our relative entropy perspective, which does not depend on Lorentz symmetry, to produce more diverse physics results.
@@ Acknowledgements
We are grateful to Toshifumi Noumi for his helpful comments and the opportunity to initiate this work. DU would like to thank the CERN Theory Department for their financial support and hospitality during this work. DU is supposed by grants from the ISF (No. 1002/23 and 597/24) and the BSF (No. 2021800).
@@ Appendix A Field redefinition of heavy fields
Appendix B Perturbative calculation of $\widetilde{\phi}_{1,\lambda}$ in eq. (28)
Let us consider $\lambda$ dependencies in $W_{\lambda}[\widetilde{\phi}_{1,\lambda}]$ and $\widetilde{\phi}_{1,\lambda}$. The classical solution $\widetilde{\phi}_{1,\lambda}$ is defined so that $\delta W_{\lambda}[\phi]/\delta\phi|_{\phi=\widetilde{\phi}_{1,\lambda}}=0$, where $W_{\lambda}[\phi]$ is given, up to the third order of $\lambda$, as
$W_{\lambda}[\phi]$ $=\int(d^{4}x)_{\rm E}\Bigg{[}a_{0}\left(\phi\right)-\lambda^{2}\frac{1}{2m_{2}^{2}\left(1-\partial^{2}_{I}/m_{2}^{2}\right)}\left(a_{1}\left(\phi\right)\right)^{2}+\lambda^{3}\frac{1}{m_{2}^{6}\left(1-\partial^{2}_{I}/m_{2}^{2}\right)^{3}}$
×$\times\bigg{\{}-a_{3}\left(0\right)\left(a_{1}\left(\phi\right)\right)^{3}+m_{2}^{2}\left(1-\frac{\partial^{2}_{I}}{m_{2}^{2}}\right)\left(a_{1}\left(\phi_{1}\right)\right)^{2}\left(a_{2}\left(\phi\right)-\frac{m_{2}^{2}}{2}\right)\bigg{\}}\Bigg{]}.$ (131)
For ease of considerations, from eq. (131), we introduce two independent parameters $\lambda^{(2)}$ and $\lambda^{(3)}$, and define as,
$W[\phi;\lambda^{(2)},\lambda^{(3)}]:$ $=\int(d^{4}x)_{\rm E}\Bigg{[}a_{0}\left(\phi\right)-\lambda^{(2)}\frac{1}{2m_{2}^{2}\left(1-\partial^{2}_{I}/m_{2}^{2}\right)}\left(a_{1}\left(\phi\right)\right)^{2}+\lambda^{(3)}\frac{1}{m_{2}^{6}\left(1-\partial^{2}_{I}/m_{2}^{2}\right)^{3}}$
×$\times\bigg{\{}-a_{3}\left(0\right)\left(a_{1}\left(\phi\right)\right)^{3}+m_{2}^{2}\left(1-\frac{\partial^{2}_{I}}{m_{2}^{2}}\right)\left(a_{1}\left(\phi_{1}\right)\right)^{2}\left(a_{2}\left(\phi\right)-\frac{m_{2}^{2}}{2}\right)\bigg{\}}\Bigg{]}.$ (132)
Regarding the two parameters $\lambda^{(2)}$ and $\lambda^{(3)}$ as expansion parameters, the solution of $W[\phi;\lambda^{(2)},\lambda^{(3)}]$ can be expressed as,
$\widetilde{\phi}_{1,\lambda}=\widetilde{\phi}_{1,0}+\delta\phi_{\lambda},$ (133)
where $\widetilde{\phi}_{1,0}$ is the classical solution of $W_{0}[\phi]=W[\phi;0,0]=\int(d^{4}x)_{\rm E}a_{0}\left(\phi\right)$, and $\delta\phi_{\lambda}$ denotes a perturbative effect coming from $\lambda^{(2)}$ and $\lambda^{(3)}$ and is expressed as a power expansion of $\lambda^{(2)}$ and $\lambda^{(3)}$. Since we can obtain the classical solution of eq. (131) from eq. (133) by replacing as $\lambda^{(2)}\to\lambda^{2}$ and $\lambda^{(3)}\to\lambda^{3}$, we find $\delta\phi_{\lambda}=\mathcal{O}\left(\lambda^{2}\right)$.
Next, let us substitute eq. (133) into eq. (131), and then we obtain as
$W_{\lambda}[\widetilde{\phi}_{1,\lambda}]$ $=\int(d^{4}x)_{\rm E}\Bigg{[}a_{0}\left(\widetilde{\phi}_{1,\lambda}\right)-\lambda^{2}\frac{1}{2m_{2}^{2}\left(1-\partial^{2}_{I}/m_{2}^{2}\right)}\left(a_{1}\left(\widetilde{\phi}_{1,\lambda}\right)\right)^{2}+\lambda^{3}\frac{1}{m_{2}^{6}\left(1-\partial^{2}_{I}/m_{2}^{2}\right)^{3}}$
×$\times\bigg{\{}-a_{3}\left(0\right)\left(a_{1}\left(\widetilde{\phi}_{1,\lambda}\right)\right)^{3}+m_{2}^{2}\left(1-\frac{\partial^{2}_{I}}{m_{2}^{2}}\right)\left(a_{1}\left(\widetilde{\phi}_{1,\lambda}\right)\right)^{2}\left(a_{2}\left(\widetilde{\phi}_{1,\lambda}\right)-\frac{m_{2}^{2}}{2}\right)\bigg{\}}\Bigg{]},$ (134)
where three functions $a_{0}\left(\widetilde{\phi}_{1,\lambda}\right)$, $a_{1}\left(\widetilde{\phi}_{1,\lambda}\right)$, and $a_{2}\left(\widetilde{\phi}_{1,\lambda}\right)$ can be expressed as
$a_{0}\left(\widetilde{\phi}_{1,\lambda}\right)$ $=a_{0}\left(\widetilde{\phi}_{1,0}\right)+\frac{\delta a_{0}}{\delta\phi}\bigg{|}_{\phi=\widetilde{\phi}_{1,0}}\delta\phi_{\lambda}+\mathcal{O}\left(\delta\phi_{\lambda}^{2}\right)=a_{0}\left(\widetilde{\phi}_{1,0}\right)+\mathcal{O}\left(\delta\phi_{\lambda}^{2}\right),$ (135)
$a_{1}\left(\widetilde{\phi}_{1,\lambda}\right)$ $=a_{1}\left(\widetilde{\phi}_{1,0}\right)+\frac{\delta a_{1}}{\delta\phi}\bigg{|}_{\phi=\widetilde{\phi}_{1,0}}\delta\phi_{\lambda}+\mathcal{O}\left(\delta\phi_{\lambda}^{2}\right),$ (136)
$a_{2}\left(\widetilde{\phi}_{1,\lambda}\right)$ $=a_{2}\left(\widetilde{\phi}_{1,0}\right)+\frac{\delta a_{2}}{\delta\phi}\bigg{|}_{\phi=\widetilde{\phi}_{1,0}}\delta\phi_{\lambda}+\mathcal{O}\left(\delta\phi_{\lambda}^{2}\right),$ (137)
where $\delta a_{0}/\delta\phi|_{\phi=\widetilde{\phi}_{1,0}}=0$ holds because $\widetilde{\phi}_{1,0}$ is the classical solution of $W_{0}[\phi]=\int(d^{4}x)_{\rm E}a_{0}\left(\phi\right)$. Note here that $\mathcal{O}\left(\delta\phi_{\lambda}^{2}\right)=\mathcal{O}\left(\lambda^{4}\right)$ because of $\delta\phi_{\lambda}=\mathcal{O}\left(\lambda^{2}\right)$. Combining eqs. (135)-(137) with eq. (134), we end up with
$W_{\lambda}[\widetilde{\phi}_{1,\lambda}]$ $=\int(d^{4}x)_{\rm E}\Bigg{[}a_{0}\left(\widetilde{\phi}_{1,0}\right)-\lambda^{2}\frac{1}{2m_{2}^{2}\left(1-\partial^{2}_{I}/m_{2}^{2}\right)}\left(a_{1}\left(\widetilde{\phi}_{1,0}\right)\right)^{2}+\lambda^{3}\frac{1}{m_{2}^{6}\left(1-\partial^{2}_{I}/m_{2}^{2}\right)^{3}}$
×$\times\bigg{\{}-a_{3}\left(0\right)\left(a_{1}\left(\widetilde{\phi}_{1,0}\right)\right)^{3}+m_{2}^{2}\left(1-\frac{\partial^{2}_{I}}{m_{2}^{2}}\right)\left(a_{1}\left(\widetilde{\phi}_{1,0}\right)\right)^{2}\left(a_{2}\left(\widetilde{\phi}_{1,0}\right)-\frac{m_{2}^{2}}{2}\right)\bigg{\}}\Bigg{]}+\mathcal{O}\left(\lambda^{4}\right).$ (138)
Appendix C Higher order of perturbative expansion in the effective action
@@
    U(1) symmetric potential in subsubsection 3.1.1 — Upon substituting eq. (77) into eq. (144), we obtain the Euclidean effective actions, up to the sixth order of $\lambda$, as
@@
    A modified version of potential in eq. (75) — Upon substituting eq. (87) into eq. (144), we obtain the Euclidean effective actions, up to the sixth order of $\lambda$, as
@@
Appendix D Derivation of eq. (48)
Returning to the definition of $\left({\partial W_{\lambda}^{(N)}}/{\partial\lambda}\right)_{\lambda=0}$ helps us understand how to calculate eq. (48). The left-hand side of eq. (48) is given by
$\left(\frac{\partial W_{\lambda}^{(N)}}{\partial\lambda}\right)_{\lambda=0}$ $=-\left(\frac{\partial\ln Z_{\lambda}^{(N)}[\widetilde{\phi}_{1,\lambda}^{(N)},\cdots,\widetilde{\phi}^{(N)}_{N,\lambda}]}{\partial\lambda}\right)_{\lambda=0}$ (156)
$=-\left(\frac{1}{Z_{\lambda}^{(N)}[\widetilde{\phi}_{1,\lambda}^{(N)},\cdots,\widetilde{\phi}^{(N)}_{N,\lambda}]}\frac{\partial}{\partial\lambda}Z_{\lambda}^{(N)}[\widetilde{\phi}_{1,\lambda}^{(N)},\cdots,\widetilde{\phi}^{(N)}_{N,\lambda}]\right)_{\lambda=0}$ (157)
$=-\left(\frac{1}{Z_{\lambda}^{(N)}[\widetilde{\phi}_{1,\lambda}^{(N)},\cdots,\widetilde{\phi}^{(N)}_{N,\lambda}]}\frac{\partial}{\partial\lambda}\int\prod_{i=1}^{N}d[\phi_{i}]e^{-I^{(N)}_{\lambda}[\phi_{1},\cdots,\phi_{N}]}\right)_{\lambda=0}$ (158)
$=\left(\frac{1}{Z_{\lambda}^{(N)}[\widetilde{\phi}_{1,\lambda}^{(N)},\cdots,\widetilde{\phi}^{(N)}_{N,\lambda}]}\int\prod_{i=1}^{N}d[\phi_{i}]e^{-I^{(N)}_{\lambda}[\phi_{1},\cdots,\phi_{N}]}\frac{\partial}{\partial\lambda}I_{\lambda}^{(N)}[\phi_{1},\cdots,\phi_{N}]\right)_{\lambda=0}$ (159)
$=\int\prod_{i=1}^{N}d[\phi_{i}]P_{\rm R}[\phi_{1},\cdots,\phi_{N}]I_{\rm I}[\phi_{1},\cdots,\phi_{N}],$ (160)
where $W_{\lambda}^{(N)}:=-\ln Z_{\lambda}^{(N)}[\widetilde{\phi}_{1,\lambda}^{(N)},\cdots,\widetilde{\phi}^{(N)}_{N,\lambda}]$, and the following relations,
$\frac{\partial}{\partial\lambda}e^{-I^{(N)}_{\lambda}[\phi_{1},\cdots,\phi_{N}]}=-e^{-I^{(N)}_{\lambda}[\phi_{1},\cdots,\phi_{N}]}\frac{\partial}{\partial\lambda}I_{\lambda}[\phi_{1},\cdots,\phi_{N}]=-I_{\rm I}[\phi_{1},\cdots,\phi_{N}]e^{-I^{(N)}_{\lambda}[\phi_{1},\cdots,\phi_{N}]},$ (161)
$P_{\rm R}[\phi_{1},\cdots,\phi_{N}]=e^{-I^{(N)}_{0}[\phi_{1},\cdots,\phi_{N}]}/Z_{0}^{(N)}[\widetilde{\phi}_{1,0}^{(N)},\cdots,\widetilde{\phi}^{(N)}_{N,0}],$ (162)
$Z_{\lambda}^{(N)}[\widetilde{\phi}_{1,\lambda}^{(N)},\cdots,\widetilde{\phi}^{(N)}_{N,\lambda}]:=\int\prod_{i=1}^{N}d[\phi_{i}]e^{-I^{(N)}_{\lambda}[\phi_{1},\cdots,\phi_{N}]},$ (163)
are used. Equation (160) was used to derive eq. (7), i.e., the right-hand side of eq. (160) is the definition of $\left({\partial W_{\lambda}^{(N)}}/{\partial\lambda}\right)_{\lambda=0}$ in this work. At the tree-level (corresponding to saddle point approximation), eq. (160) is calculated as follows:
$\left(\frac{\partial W_{\lambda}^{(N)}}{\partial\lambda}\right)_{\lambda=0}$ $=\int\prod_{i=1}^{N}d[\phi_{i}]\frac{1}{Z_{0}^{(N)}[\widetilde{\phi}_{1,0}^{(N)},\cdots,\widetilde{\phi}^{(N)}_{N,0}]}e^{-I^{(N)}_{0}[\phi_{1},\cdots,\phi_{N}]}I_{\rm I}[\phi_{1},\cdots,\phi_{N}]$
$=\frac{1}{Z_{0}^{(N)}[\widetilde{\phi}_{1,0}^{(N)},\cdots,\widetilde{\phi}^{(N)}_{N,0}]}e^{-I^{(N)}_{0}[\widetilde{\phi}_{1,0}^{(N)},\cdots,\widetilde{\phi}^{(N)}_{N,0}]}I_{\rm I}[\widetilde{\phi}_{1,0}^{(N)},\cdots,\widetilde{\phi}^{(N)}_{N,0}]$
$=I_{\rm I}[\widetilde{\phi}_{1,0}^{(N)},\cdots,\widetilde{\phi}^{(N)}_{N,0}],$ (164)
where we used the following relation:
$Z_{0}^{(N)}[\widetilde{\phi}_{1,0}^{(N)},\cdots,\widetilde{\phi}^{(N)}_{N,0}]=\int\prod_{i=1}^{N}d[\phi_{i}]e^{-I^{(N)}_{0}[\phi_{1},\cdots,\phi_{N}]}=e^{-I^{(N)}_{0}[\widetilde{\phi}_{1,0}^{(N)},\cdots,\widetilde{\phi}_{N,0}^{(N)}]},$ (165)
with the saddle point approximation. From eq. (40) and the solution $\widetilde{\phi}^{(N)}_{N,0}=0$, we find $I_{\rm I}[\widetilde{\phi}_{1,0}^{(N)},\cdots,\widetilde{\phi}^{(N)}_{N,0}]=0$. And then, from eq. (164), we end up with
$\left(\frac{\partial W^{(N)}_{\lambda}}{\partial\lambda}\right)_{\lambda=0}=0.$ (166)
Appendix E Relation between mass bases and interaction bases of the fields